OpenStudy (anonymous):

how would you get rid of the fraction within a fraction again? I forgot :P how would you get rid of the fraction within a fraction again? I forgot :P @Mathematics

OpenStudy (anonymous):

Let me give an example

OpenStudy (anonymous):

\[(2/\sqrt{3})\div (\sqrt{3}/2)\]

OpenStudy (zarkon):

\[\frac{2}{\sqrt{3}}\div \frac{\sqrt{3}}{2}\] \[=\frac{2}{\sqrt{3}}\times \frac{2}{\sqrt{3}}\]

OpenStudy (anonymous):

So it would be 4/2sqrt(3)???

OpenStudy (zarkon):

\[\frac{4}{3}\]